Check my Exercise 5-17 Break-Even and Target Profit Analysis [LO5-4, LO5-5, LO5-6] Outback Outfitters sells recreational equiRequired 1 Required 2 Required 3 Required 4 What is the break-even point in unit sales and in dollar sales? Break-even point2.85 points Complete this question by entering your answers in the tabs below. eBook Required 1 Required 2 Required 3 RequireAt present, the company is selling 15,000 stoves per month. The sales manager is convinced that a 10% reduction in the sellinComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Refer to the d
Answer #1

Answers

  • [1]

A

Sale price

$140

B

Variable Cost

$98

C = A - B

Unit contribution margin

$42

D

Fixed Cost

$184,800

E = D/C

Break even point in units

4400

Answer

F = E x A

Break even point in $

$616,000

Answer

  • [2]
    Higher Break even point will arise, because higher variable cost % = lower CM ratio = higher break even point.
  • [3]

Present

Proposed

15000

stoves

18750

stoves

Total

per unit

Total

per unit

Sale

$2,100,000

$140

$2,362,500

$126

Variable cost

$1,470,000

$98

$1,837,500

$98

Contribution margin

$630,000

$42

$525,000

$28

Fixed cost

$184,800

$184,800

Net Income or net operating income

$445,200

$340,200

  • [4]

A

New Sale price

$                  126.00

B

Variable cost

$                    98.00

C = A - B

New Contribution margin per unit

$                    28.00

D

Fixed Cost

$          184,800.00

E

Target Profits

$            75,000.00

F = D+E

Total Contribution margin required

$          259,800.00

G = F/C

No. of stoves to be sold to earn target Profits

                        9,279 ANSWER